I need help solving this word problem I tried so many time and I do not get the right answer..thank you in advance
In planning for a new item, a manufacturer assumes that the number of items produced x and the cost in dollars C of producing these items are related by a linear equation. Projections are that 100 items will cost $10,000 to produce and that 300 items will cost $22,000 to produce. how can i Find the equation that relates C and x.

Answer by stanbon(48558)

In planning for a new item, a manufacturer assumes that the number of items produced x and the cost in dollars C of producing these items are related by a linear equation. Projections are that 100 items will cost $10,000 to produce and that 300 items will cost $22,000 to produce. how can i Find the equation that relates C and x.
------------
You have two (x,C) points: (100,10000) and (300,22000)
Find the slope and y-intercept so you can find the equation.
----------
slope = [22000-10000]/[300-100] = 12000/200 = 60
--------
10000 = 60(100)+b
10000 = 6000 + b
b= 4000
Cost = 60(#of items) + 4000
C = 60x + 4000
============
Cheers,
Stan H.